OpenStudy (anonymous):

Find the solution to the system of equations by using either graphing or substitution. y = 6 – x and y = x – 2 (1 point) (2, 4) (–4, 2) (4, 2) no solutions

OpenStudy (mendicant_bias):

Using substitution: Given the equations that you have, \[y = 6 -x\]\[y = x -2\]What you need to do is isolate one of the variables, such a y, in one of the equations and plug it as a placeholder like such: \[y = 6 - x\]Therefore,\[6 - x = x - 2\]From here, you just need to simplify. Then you can plug back in the value you got for whatever variable you solved for. Make sense?

OpenStudy (mendicant_bias):

Does this help?

OpenStudy (anonymous):

yeah! is it C